OpenStudy (anonymous):

Simplify. 5x / x - 3 + 6x / x - 3

OpenStudy (anonymous):

\[\frac{ 5x }{ x - 3 } + \frac{ 6x }{ x - 3 }\]

OpenStudy (anonymous):

Can your answer be a decimal? Because I presume you add the like terms. So 5 + 6 = 11 and 3 + 3 = 6. Then divide 11 into 6 and the answer, 1.83 That if your answer can be a decimal, otherwise I don't remember how to do these. Sorry about that!